# TrustGeoGen: Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

### v2

@@ 3 Methodology
3.1 Data Engine
@@ 10 $s\leftarrow S_{i-1}$;
12
13Assess path length: $L\leftarrow|\mathcal{P}|$;
@@ 20 Discard $\mathcal{P}$;
$\small\mathcal{G}=(S,S_{0},R,\hookrightarrow),$ (1)
$\small S=S_{0}\cup\left\{s^{\prime}\mid\forall S_{r}\subset S,\,r\in R,\,S_{r}\xhookrightarrow{r}s^{\prime}\right\}.$ (2)
@@
$\small\mathcal{P}=\{(S_{i-1},r_{s},s)\mid\forall s\in S_{i},S_{i-1}\xhookrightarrow{r_{s}}s,i=n,\dots,1\},$ (3)
@@ 30    Break;
31
32
33return $\mathcal{P}_{\text{set}}$;
@@ 10   Add $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ to $\mathcal{D}$;
3.4 Self-reflective Traceback Data
@@
Recent studies on the reasoning capabilities of large models suggest that these models acquire substantial knowledge during the pre-training phase. By providing training data with various cognitive templates during the post-training phase, the models can better utilize their inherent knowledge, thereby enhancing reasoning ability and improving response quality(ye2025limo, ; yang2025reasonflux, ). Among these cognitive templates, self-reflection trace-back reasoning serve as highly effective approaches. However, self-reflection and retrospective reasoning often occur during the author’s cognitive process and are rarely preserved in the final outputs, making the collection of relevant data particularly challenging. These types of data are even more scarce in the domain of multi-modal geometric problem-solving.
@@ 4.1 Data Distribution
LLaVA-1.5-7B (llava/liu2024visual, ) 7B 4 1.67% 3 5.00% 1 0.42% 0 0.00% 0 0.00%
Qwen2-VL-7B (Qwen2-VL, ) 7B 11 4.58% 5 8.33% 2 3.33% 2 3.33% 2 3.33%
GPT-4o openai2024gpt4o - 62 25.83% 31 51.67% 10 16.67% 11 18.33% 10 16.67%
Claude-3.7-sonnet (2024claude, ) - 66 27.50% 33 55.00% 16 26.67% 10 16.67% 7 11.67%
Qwen2.5-VL-72B Qwen2.5-VL 72B 68 28.33% 32 53.33% 15 25.00% 15 25.00% 6 10.00%
DeepSeek-R1 * (DeepSeekAI2025DeepSeekR1IR, ) 671B 89 37.08% 30 50.00% 20 33.33% 22 36.67% 17 28.33%
Intern-S1 bai2025intern 235B+6B 104 43.33% 35 58.33% 21 35.00% 25 41.67% 23 38.33%
Gemini-2.5-pro comanici2025gemini - 104 43.33% 34 56.67% 24 40.00% 26 43.33% 20 33.33%
OpenAI-o3 openai_o3 - 110 45.83% 37 61.67% 24 40.00% 26 43.33% 23 38.33%
@@
4.2 Deep Reasoning Augmentation
@@
4.3 Construction and Analysis of GeoTrust-Test
@@ 5 Experiments
  • Whether existing MLLMs demonstrate competent performance on complex geometric problems?
  • Whether the data generated by TrustGeoGen have sufficient complexity to provide gains for SOTA models?
  • Whether geometric reasoning data with rigorous formal verification improve MLLMs’ capabilities and provide specific advantages?
  • Whether GeoTrust, constructed without prior data sources, can generalize to OOD geometric testset?
@@
Model Release Date
GeoQA
(mid-school level)
Geometry3K
(mid-school level)
OlympiadBench-Geo
(olympiad level)
GeoTrust-test
GPT-4o openai2024gpt4o May, 2024 42.31% 31.45% 13.39% 25.83%
Claude-3.7-sonnet (2024claude, ) Feb, 2025 49.73% 33.28% 17.86% 27.50%
Qwen2.5-VL-72B Qwen2.5-VL Jan, 2025 67.90% 35.44% 29.46% 28.33%
Intern-S1 bai2025intern Jul, 2025 62.47% 52.31% 49.11% 43.33%
Gemini-2.5-procomanici2025gemini Jun, 2025 79.58% 80.70% 75.00% 43.33%
OpenAI-o3 openai_o3 Apr, 2025 79.31% 81.03% 77.68% 45.83%
5.1 Dataset, Metric, and Implementation Detail
Dataset. To evaluate the capability of MLLMs in solving complex geometry problems, we construct GeoTrust-test by partitioning the original GeoTrust dataset. As introduced in section˜4.3, the testset GeoTrust-test comprehensively covers four difficulty levels ranging from $Tier_{1}$ to $Tier_{4}$. To validate the effectiveness of our data construction, we additionally sample 2342 geometry problems as GeoTrust-train, with different difficulty tier.
@@ 5.1 Dataset, Metric, and Implementation Detail
5.2 Performance of Existing MLLMs
@@
5.3 Complexity Analysis Between Benchmarks
@@
5.4 Effectiveness Analysis of Trustworthy Data
TrustGeoGen constructs trustworthy geometric reasoning data using a formal engine. It bridges the gap between formal and human-like reasoning through "connection thinking" (algorithm˜1). Furthermore, thinking template data—comprising multi-solution data (section˜3.3) and self-reflective traceback data (section˜3.4)—is utilized to introduce diverse reasoning pathways. In this section, we design comprehensive experiments to validate the effectiveness of our trustworthy data.
@@ 5.4 Effectiveness Analysis of Trustworthy Data
5.4.1 Necessary of Connection Thinking
@@
5.4.2 Advantages Provided by Trustworthy Data
To validate the effectiveness of data generated by TrustGeoGen, we conducted a comparative experiment using pseudo-labeled data. We fed 2,158 questions generated by TrustGeoGen into OpenAI-o3 openai_o3 and obtained answers from the model. The responses were then evaluated using the metrics described in section˜5.1, resulting in 846 correctly answered questions, which we treat as pseudo-labeled data. For fairness, we also used the same 846 questions as the GeoTrust Data in our experiments.
@@ 5.4.2 Advantages Provided by Trustworthy Data
5.4.3 Advantages of by Thinking Template
@@ 5.5.1 GeoQA
GeoQA was evaluated in a direct question-answering (QA) format, rather than a multiple-choice setting. The same metrics as described in section˜5.1 were used.
5.5.2 Geometry3K
@@
Geometry3K was evaluated in a direct QA format, rather than a multiple-choice setting. The same metrics as described in section˜5.1 were used.
5.5.3 OlympiadBench-geo
@@
5.5.4 Ablation Study of Thinking Template
As established in the preceding sections, our GeoTrust-train dataset is composed of two primary components: base data and thinking template (TT) data. To validate the effectiveness of the thinking template data, we conduct a series of ablation studies on the GeoQA and Geometry3K datasets. Following the same methodology outlined in section˜5.4.3, we partition the GeoTrust-train data into three distinct configurations for our experiments: without thinking templates (w/o TT), replace thinking templates (Re TT), and full data. Each configuration is then combined with the training sets of GeoQA and Geometry3K, respectively, to fine-tune the models. The models are subsequently evaluated on the corresponding test sets.
@@ 5.5.4 Ablation Study of Thinking Template
5.5.5 Summary of Experiments Results
@@
6 Conclusion and Further Discussion

### v3

@@ III Methodology
Figure 1: Overview of TrustGeoGen: (1) Constructor builds geometric premises and visual diagrams; (2) Reasoner generates a formally validated reasoning graph; (3) Sampler utilizes GeoExplore algorithms to select conclusions and trace diverse reasoning paths; (4) Translator converts formal logic into fluent natural language using “connection thinking”. A bootstrap mechanism amplifies problem difficulty iteratively.
III-A Data Engine
@@
Figure 2: Simple illustration of how Translator works
$\mathcal{G}=(S,S_{0},R,\hookrightarrow),$ (1)
$S=S_{0}\cup\left\{s^{\prime}\mid\forall S_{r}\subset S,\,r\in R,\,S_{r}\xhookrightarrow{r}s^{\prime}\right\}.$ (2)
@@ 10 $s\leftarrow S_{i-1}$;
12 end while
13
14Assess path length: $L\leftarrow|\mathcal{P}|$;
@@ 22 Discard $\mathcal{P}$;
$\mathcal{P}=\{(S_{i-1},r_{s},s)\mid\forall s\in S_{i},S_{i-1}\xhookrightarrow{r_{s}}s,i=n,\dots,1\},$ (3)
@@ 33    Break;
34
35   end if
36
37 end while
38return $\mathcal{P}_{\text{set}}$;
@@ III-C Multi-solution Data
III-D Self-reflective Traceback Data
@@
Recent studies on the reasoning capabilities of large models suggest that these models acquire substantial knowledge during the pre-training phase. By providing training data with various cognitive templates during the post-training phase, the models can better utilize their inherent knowledge, thereby enhancing reasoning ability and improving response quality[ye2025limo, yang2025reasonflux]. Among these cognitive templates, self-reflection trace-back reasoning serve as highly effective approaches. However, self-reflection and retrospective reasoning often occur during the author’s cognitive process and are rarely preserved in the final outputs, making the collection of relevant data particularly challenging. These types of data are even more scarce in the domain of multi-modal geometric problem-solving.
@@ IV-A Data Distribution
LLaVA-1.5-7B [llava/liu2024visual] 7B 4 1.67% 3 5.00% 1 0.42% 0 0.00% 0 0.00%
Qwen2-VL-7B [Qwen2-VL] 7B 11 4.58% 5 8.33% 2 3.33% 2 3.33% 2 3.33%
GPT-4o [openai2024gpt4o] - 62 25.83% 31 51.67% 10 16.67% 11 18.33% 10 16.67%
Claude-3.7-sonnet [2024claude] - 66 27.50% 33 55.00% 16 26.67% 10 16.67% 7 11.67%
Qwen2.5-VL-72B [Qwen2.5-VL] 72B 68 28.33% 32 53.33% 15 25.00% 15 25.00% 6 10.00%
DeepSeek-R1 * [DeepSeekAI2025DeepSeekR1IR] 671B 89 37.08% 30 50.00% 20 33.33% 22 36.67% 17 28.33%
Intern-S1 [bai2025intern] 235B+6B 104 43.33% 35 58.33% 21 35.00% 25 41.67% 23 38.33%
Gemini-2.5-pro [comanici2025gemini] - 104 43.33% 34 56.67% 24 40.00% 26 43.33% 20 33.33%
OpenAI-o3 [openai_o3] - 110 45.83% 37 61.67% 24 40.00% 26 43.33% 23 38.33%
@@
IV-B Deep Reasoning Augmentation
@@
IV-C Construction and Analysis of GeoTrust-Test
@@ V Experiments
  • Whether existing MLLMs demonstrate competent performance on complex geometric problems? (Section˜V-B)
  • Whether the data generated by TrustGeoGen have sufficient complexity to provide gains for SOTA models? (Section˜V-C)
  • Whether geometric reasoning data with rigorous formal verification improve MLLMs’ capabilities and provide specific advantages? (Section˜V-D)
  • Whether GeoTrust, constructed without prior data sources, can generalize to OOD geometric testset? (Section˜V-E)
@@
Model Release Date GeoQA (mid-school level) Geometry3K (mid-school level) OlympiadBench-Geo (olympiad level) GeoTrust-test
GPT-4o [openai2024gpt4o] May, 2024 42.31% 31.45% 13.39% 25.83%
Claude-3.7-sonnet [2024claude] Feb, 2025 49.73% 33.28% 17.86% 27.50%
Qwen2.5-VL-72B [Qwen2.5-VL] Jan, 2025 67.90% 35.44% 29.46% 28.33%
Intern-S1 [bai2025intern] Jul, 2025 62.47% 52.31% 49.11% 43.33%
Gemini-2.5-pro[comanici2025gemini] Jun, 2025 79.58% 80.70% 75.00% 43.33%
OpenAI-o3 [openai_o3] Apr, 2025 79.31% 81.03% 77.68% 45.83%
V-A Dataset, Metric, and Implementation Detail
Dataset. To evaluate the capability of MLLMs in solving complex geometry problems, we construct GeoTrust-test by partitioning the original GeoTrust dataset. As introduced in Section˜IV-C, the testset GeoTrust-test comprehensively covers four difficulty levels ranging from $Tier_{1}$ to $Tier_{4}$. To validate the effectiveness of our data construction, we additionally sample 2342 geometry problems as GeoTrust-train, with different difficulty tier.
@@ V-A Dataset, Metric, and Implementation Detail
V-B Performance of Existing MLLMs
@@
V-C Complexity Analysis Between Benchmarks
@@
V-D Effectiveness Analysis of Trustworthy Data
TrustGeoGen constructs trustworthy geometric reasoning data using a formal engine. It bridges the gap between formal and human-like reasoning through "connection thinking" (Algorithm˜2). Furthermore, thinking template data—comprising multi-solution data (Section˜III-C) and self-reflective traceback data (Section˜III-D)—is utilized to introduce diverse reasoning pathways. In this section, we design comprehensive experiments to validate the effectiveness of our trustworthy data.
@@ V-D Effectiveness Analysis of Trustworthy Data
V-D1 Necessary of Connection Thinking
@@
V-D2 Advantages Provided by Trustworthy Data
To validate the effectiveness of data generated by TrustGeoGen, we conducted a comparative experiment using pseudo-labeled data. We fed 2,158 questions generated by TrustGeoGen into OpenAI-o3 [openai_o3] and obtained answers from the model. The responses were then evaluated using the metrics described in Section˜V-A, resulting in 846 correctly answered questions, which we treat as pseudo-labeled data. For fairness, we also used the same 846 questions as the GeoTrust Data in our experiments.
@@ V-D2 Advantages Provided by Trustworthy Data
V-D3 Advantages of by Thinking Template
@@ V-E1 GeoQA
GeoQA was evaluated in a direct question-answering (QA) format, rather than a multiple-choice setting. The same metrics as described in Section˜V-A were used.
V-E2 Geometry3K
@@
Geometry3K was evaluated in a direct QA format, rather than a multiple-choice setting. The same metrics as described in Section˜V-A were used.
V-E3 OlympiadBench-geo
@@
V-E4 Ablation Study of Thinking Template
As established in the preceding sections, our GeoTrust-train dataset is composed of two primary components: base data and thinking template (TT) data. To validate the effectiveness of the thinking template data, we conduct a series of ablation studies on the GeoQA and Geometry3K datasets. Following the same methodology outlined in Section˜V-D3, we partition the GeoTrust-train data into three distinct configurations for our experiments: without thinking templates (w/o TT), replace thinking templates (Re TT), and full data. Each configuration is then combined with the training sets of GeoQA and Geometry3K, respectively, to fine-tune the models. The models are subsequently evaluated on the corresponding test sets.
@@ V-E4 Ablation Study of Thinking Template
V-E5 Summary of Experiments Results
@@
VI Conclusion and Further Discussion